OpenStudy (anonymous):

The "line of best fit" for a scatterplot must pass through at least _____ of the data points in the data set. A.0 B.1 C.2 D.3

OpenStudy (amistre64):

none of them really; its a sort of mean or average; and those dont need to be apart of the set

OpenStudy (amistre64):

th eline of best fit MUST use the average X and Y values in its construction; but that average maynot necessarily be any of the stated points

OpenStudy (amistre64):

for example: x y xy 0 0 0 1 2 2 2 5 10 --- 12 sum x sumx = 9 sum y sumy = 49 sumx sumy = 21 avgX = 1 avgY = 7/3 m = 3(12) - 21 --------- = 3/2 3(5) - 9 y = 3/2 (x-1) + 7/3 should be the line of best fit for those 3 points

OpenStudy (anonymous):

oh okay thanks c:

OpenStudy (amistre64):

yep

OpenStudy (anonymous):

Tell whether the lines for each pair of equations are parallel, perpendicular, or neither. y = 5x y = -5x + 7 a. parallel b. neither c. perpendicular

OpenStudy (anonymous):

can u help